SpaceRef

SpaceRef


NOAA SATOPS Morning Report: Wednesday, May 23, 2007

Status Report From: NOAA Office of Satellite Operations
Posted: Wednesday, May 23, 2007

Customers or users who want to call to ask about problems or report outages can use the following contacts:

- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444

- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

J143

NOTE: NOAA-14 is scheduled to be decommissioned today (5/23/2007) at 10:44 EDT.

GOES

All GOES operations were nominal over the past 24 hours. No anomalies affecting product delivery or s/c health & safety. For GOES status and other information please see: http://www.oso.noaa.gov/goes/index.htm

POES

NOAA-18; Rev 10326; 1730Z at Wallops: 27.1 percent of HRPT data was lost due to the prime antenna losing track. Wallops regained track with the backup antenna. The operations team was able to fully recover all LAC data.

All POES operations were nominal over the past 24 hours. No anomalies affecting product delivery or s/c health & safety. For a detailed status and description of POES see: http://www.oso.noaa.gov/poes/index.htm

DMSP

All DMSP operations were nominal over the past 24 hours. No nomalies affecting product delivery or s/c health & safety. For more DMSP Information please see: http://dmsp.ngdc.noaa.gov/dmsp.html

Questions or comments to Ron.Mahmot@noaa.gov

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter